# Two-Stage Nuisance Function Estimation for Causal Mediation Analysis

### v1

@@ Abstract
When estimating the direct and indirect causal effects using the influence function-based estimator of the mediation functional, it is crucial to understand what aspects of the treatment, the mediator, and the outcome mean mechanisms should be focused on. Specifically, considering them as nuisance functions and attempting to fit these nuisance functions as accurate as possible is not necessarily the best approach to take. In this work, we propose a two-stage estimation strategy for the nuisance functions that estimates the nuisance functions based on the role they play in the structure of the bias of the influence function-based estimator of the mediation functional. We provide robustness analysis of the proposed method, as well as sufficient conditions for consistency and asymptotic normality of the estimator of the parameter of interest.
@@ 1 Introduction
The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model for the data generating process, the direct and indirect effects can be identified by a functional known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those assumptions hold in the system under study and the parameter is identified, the next question is how to estimate the mediation functional.
@@
There has been several estimation strategies proposed in the literature based on regression and weighting techniques. In their seminal paper, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric effeicint and has the desired property that under certain regularity conditions, it remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism, one describing the mediator mechanism, and one describing the outcome mean mechanism, and the model for two out of three should be correctly specified. Among the aforementioned nuisance functions, the mediator mechanism, that is, the conditional distribution of the mediator variable given covariates and the treatment, is usually a difficult object to estimate. Trying to naively fit this conditional distribution as good as possible will most likely lead to a final estimator of the mediation functional that is sensitive to even small errors in the estimator of this nuisance function. More importantly, in the process of estimation, it is not clear what aspects of this object should be focused on to obtain a better estimate of the mediation functional.
@@
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis, which avoids direct estimation of the mediator mechanism. In the first stage, we estimate two basic nuisance functions, namely the treatment mechanism and the outcome mean mechanism. In the second stage, using the basic nuisance functions, we estimate two additional nuisance functions that contain the needed information regarding the mediator mechanism. We use the weighted balancing approach of Imai and Ratkovic, (2014) to design the estimator of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide parametric and nonparametric versions of the balancing estimators. The other two nuisance functions are obtained using standard parametric or nonparametric regressions. After obtaining all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We provide a robustness analysis of the resulting estimator and demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain a consistent estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest.
@@
The rest of the paper is organized as follows. In Section 2, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. In Section 3, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. We propose our two-stage nuisance function estimation approach in Section 4, and we provide the robustness and asymptotic analyses of the proposed estimator in Section 5. We conclude in Section 6.
@@ 2 Model Description
We consider a model with binary treatment variable $A\in\{0,1\}$, observed pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$, and $Y^{(m)}=Y^{(A,m)}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $10<p(m\mid a,x)<$ 1, and $10<p(a\mid x)<$ 1.
@@
    $Y^{(a,m)}\perp\mkern-9.5mu\perp A\mid X$,
@@
$\mathbb{E}[Y^{(1,M^{(1)})}]=\mathbb{E}[Y^{(1)}]=\mathbb{E}[Y^{(1)}\mid A=1]=\mathbb{E}[Y\mid A=1].$
@@ 3 Influence Function-Based Estimator
Tchetgen Tchetgen and Shpitser, (2012) showed that the efficienet influence function of the mediation functional is given by
@@
  • $p(M\mid A=a,X)$, $\forall M\in\mathcal{M},X\in\mathcal{X}$.
@@ 4 Proposed Estimation Approach
$\pi_{1}(X)=p(A=0\mid X),$ $\forall X\in\mathcal{X},$
$\pi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M\mid A=1,X)},$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
@@
We start by presenting a high-level view of our proposed approach.
Estimating the nuisance functions: In order for estimating the nuisance functions, we proceed in the following two stages.
@@
      We choose $\hat{\pi}_{1}$ such that it satisfies
      $\mathbb{E}\left[\left\{\frac{1-A}{\hat{\pi}_{1}(X)}-1\right\}h_{1}(X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1},$ (1)
      where $\mathcal{H}_{1}\subset L^{2}(X)$ is a function space chosen by the researcher.
    • We estimate $\mu_{1}$ by regressing $Y$ on $\{M,X,A=1\}$.
@@
      Given $\hat{\pi}_{1}$ from Stage 1, we choose $\hat{\pi}_{2}$ such that it satisfies
      $\mathbb{E}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{1-A}{\hat{\pi}_{1}(X)}\right\}h_{2}(M,X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2},$ (2)
      where $\mathcal{H}_{2}\subset L^{2}(M,X)$ is a function space chosen by the researcher.
    • Given $\hat{\mu}_{1}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1}$ on $\{X,A=0\}$.
Estimating the parameter of interest: Our final estimator of the parameter of interest will be the following.
$\hat{\psi}^{mr}=\mathbb{E}_{n}\left[\hat{\mu}_{2}(X)+\frac{(1-A)}{\hat{\pi}_{1}(X)}\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{2}(X)\}+\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}\{Y-\hat{\mu}_{1}(M,X)\}\right].$
In Stage 1, we can choose to use a parametric model $\{\pi_{1}(X;\alpha):\alpha\in\Xi_{\alpha}\}$, parametrized by $\alpha$ for $\pi_{1}$. Suppose $\Xi_{\alpha}=\mathbb{R}^{d}$. Then one can choose $d$ score functions $\{h_{1,1},...,h_{1,d}\}$ from $\mathcal{H}_{1}$, and obtain $\hat{\alpha}$ using generalized method of moments (GMM) by solving the following system of $d$ equations.
$\begin{cases}&\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X;\hat{\alpha})}-1\right\}h_{1,1}(X)\right]=0\\ &\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X;\hat{\alpha})}-1\right\}h_{1,2}(X)\right]=0\\ &\vdots\\ &\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X;\hat{\alpha})}-1\right\}h_{1,d}(X)\right]=0\end{cases}$
Alternatively, one can choose to proceed with a nonparametric estimator. A naive approach would be by solving the following optimization problem.
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)\right],$
where $\Pi_{1}\subset L^{2}(X)$ is a function space chosen by the researcher. Unfortunately, this estimator will be unstable and highly sensitive to noise. The following result suggests an alternative.
@@
$\hat{\pi}_{1}$ obtained from the following minimax optimization satisfies equation (1) with equality.
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(X)}\max_{h_{1}\in L^{2}(X)}\mathbb{E}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
@@
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{\frac{1-A}{\pi_{1}(X)}-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$ (3)
@@
Note that our proposed nonparametric estimator $\hat{\pi}_{1}$ in equation (3) can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect.
Similarly, in Stage 2, to estimate $\pi_{2}$, we can use a GMM estimator, or use the following counterpart of estimator (3).
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{\frac{A\hat{\pi}_{2}(M,X)}{1-\hat{\pi}_{1}(X)}-\frac{1-A}{\hat{\pi}_{1}(X)}\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right],$ (4)
where $\Pi_{2}\subset L^{2}(M,X)$ is a function space chosen by the researcher.
@@
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, any standard parametric or nonparametric regression method can be used.
5 Analyses

### v2

@@ Abstract
Tchetgen Tchetgen and Shpitser (2012) introduced an efficient, debiased, and robust influence function-based estimator for the mediation functional, which is the key component in mediation analysis. This estimator relies on the treatment, mediator, and outcome mean mechanisms. However, treating these three mechanisms as nuisance functions and fitting them as accurately as possible may not be the most effective approach. Instead, it is essential to identify the specific functionals and aspects of these mechanisms that impact the estimation of the mediation functional. In this work, we propose a two-stage estimation strategy for certain nuisance functions in the influence function of the mediation functional that are based on these three mechanisms. This strategy is guided by the role those nuisance functions play in the bias structure of the influence function-based estimator for the mediation functional. In the first stage, we estimate two primary nuisance functions, namely the inverse treatment mechanism and the outcome mean mechanism. In the second stage, we leverage these primary functions to estimate two additional nuisance functions that encapsulate the needed information about the mediator mechanism. We propose a nonparametric weighted balancing estimation approach to design the estimator for one of the nuisance functions in Stage 1 and one of in Stage 2, where the weights are designed directly based on the bias of the final estimator for the mediation functional. The remaining two nuisance functions are estimated using standard parametric or nonparametric regression methods. Once all four nuisance functions are obtained, they are incorporated into the influence function-based estimator. We provide a robustness analysis of the proposed method and establish sufficient conditions for consistency and asymptotic normality of our estimator for the mediation functional.
@@ 1 Introduction
The idea of causal mediation analysis is to partition the total causal effect of the treatment on the outcome into a direct part and an indirect part, where the indirect part is mediated by post-treatment mediator variables. A fundamental result in the literature of causal inference demonstrates that under certain conditions on the statistical model for the data generating process, the direct and indirect effects can be identified by a functional known as the mediation formula or mediation functional (Robins and Greenland,, 1992; Pearl,, 2001). If the researcher, based on expert and background knowledge, believes those assumptions hold in the setting under study and the parameter is identified, the next question is how to estimate the mediation functional.
@@
There has been several estimation strategies proposed in the literature based on regression and weighting techniques. In their seminal paper, Tchetgen Tchetgen and Shpitser, (2012) proposed an estimator for the mediation functional which is locally semiparametric efficient and has the desired property that under certain regularity conditions, it remains consistent even if the working models for all the nuisance functions used in the estimator are not correctly specified. More specifically, they utilize one nuisance function describing the treatment mechanism (i.e., the conditional distribution of the treatment given the covariates), one describing the mediator mechanism (i.e., the conditional distribution of the mediator variable given covariates and the treatment), and one describing the outcome mean mechanism (i.e., the conditional expectation of the outcome variable given covariates, the treatment, and the mediator), and the model for two out of three should be correctly specified. There are two challenges with their proposed estimator: (1) among the aforementioned nuisance functions, the mediator mechanism is usually a difficult object to estimate, as it may be the density of a multivariate continuous variable conditioned on another multivariate continuous variable; (2) in their estimator, the estimator of the treatment and mediator mechanisms appear in the denominator, and hence can significantly inflate the variance of the estimator, especially when these quantities are small. Therefore, trying to naively fit these conditional distributions as good as possible will not necessarily lead to an estimator with reasonable mean squared error. Instead, it is important to understand what aspects of these objects should be focused on to obtain a better estimate of the mediation functional.
@@
In this paper, we propose a two-stage nuisance function estimation approach for mediation analysis, which avoids direct estimation of the treatment and mediator mechanisms. In the first stage, we estimate two basic nuisance functions, namely the inverse of the treatment mechanism and the outcome mean mechanism. In the second stage, using the basic nuisance functions, we estimate two additional nuisance functions that contain the needed information regarding the mediator mechanism. We propose a non-parametric version of the weighted balancing approach of Imai and Ratkovic, (2014) to design the estimator of the treatment mechanism in Stage 1 and one of the nuisance functions in Stage 2. The weights in these balancing estimators are designed directly based on the bias of the final estimator of the parameter of interest. We provide an implementation of our non-parametric estimators with RKHSes as the function classes, and provide closed-form solutions for the nuisance parameters. The other two nuisance functions are obtained using standard parametric or nonparametric regressions. After obtaining all four nuisance functions, they are used in the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012). We provide a robustness analysis of the resulting estimator and demonstrate that our approach retains the multiple robustness property in the sense that not all the nuisance functions are needed to be correctly specified to obtain a asymptotically unbiased estimator of the parameter of interest. We also provide requirements on the nuisance function estimators sufficient to obtain a consistent and asymptotically normal estimator for the parameter of interest.
@@
The rest of the paper is organized as follows. In Section 2, we describe the model and review conditions sufficient for identification of the direct and indirect causal effects. In Section 3, we review the influence function-based estimator of Tchetgen Tchetgen and Shpitser, (2012) and its multiple-robustness property. We propose our two-stage nuisance function estimation approach as well as its implementation using RKHSes in Section 4. We provide the robustness and asymptotic analyses of the proposed estimator in Section 5. We conclude in Section 6.
@@ 2 Model Description
We consider a model with binary treatment variable $A\in\{0,1\}$, pre-treatment covariates $X\in\mathcal{X}$, post-treatment mediator variables $M\in\mathcal{M}$, and outcome variable $Y\in\mathcal{Y}$. Let $P_{0}$ represent the true underlying distribution from which the independent and identically distributed observations $\{O_{i}:=(X,A,M,Y)_{i}\}_{i=1}^{n}$ are collected. In this work, we use the potential outcomes framework (Rubin,, 1974). We denote the potential outcome variable of $Y$, had the treatment and mediator variables been set to value $A=a$ and $M=m$ (possibly contrary to the fact) by $Y^{(a,m)}$. Similarly, we define $M^{(a)}$ as the potential outcome variable of $M$ had the treatment variables been set to value $A=a$. Based on variables $Y^{(a,m)}$ and $M^{(a)}$, we define $Y^{(a)}=Y^{(a,M^{(a)})}$, and $Y^{(m)}=Y^{(A,m)}$. Following majority of the work in the field of causal inference, we require consistency, meaning that $M^{(a)}=M$ if $A=a$, and $Y^{(a,m)}=Y$ if $A=a$ and $M=m$, as well as positivity, meaning that for all $m$, $a$, and $x$, we have $p(m\mid a,x)>0$, and $p(a\mid x)>0$.
@@
    $Y^{(a,m)}\perp\mkern-9.5mu\perp\{A,M\}\mid X$,
@@
$\mathbb{E}[Y^{(1,M^{(1)})}]=\mathbb{E}[Y^{(1)}]=\mathbb{E}[\mathbb{E}[Y^{(1)}\mid X]]=\mathbb{E}[\mathbb{E}[Y^{(1)}\mid A=1,X]]=\mathbb{E}[\mathbb{E}[Y\mid A=1,X]].$
@@ 3 Influence Function-Based Estimator
Tchetgen Tchetgen and Shpitser, (2012) showed that in a nonparametric model, the efficient influence function of the mediation functional is given by
@@
  • $p(M\mid A=a,X)$, $\forall M\in\mathcal{M},X\in\mathcal{X},a\in\{0,1\}$.
@@
As mentioned in the introduction, there are two concerns of complexity of the nuisance functions and potential large variance about $\hat{\psi}^{TTS}$. In the next section, we propose an alternative estimation strategy that mitigates these issues.
@@ 4 Proposed Estimation Approach
$\pi_{1}(X)=\frac{1}{p(A=0\mid X)},$ $\forall X\in\mathcal{X},$
$\pi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M,A=1\mid X)},$ $\forall M\in\mathcal{M},X\in\mathcal{X},$
@@
We start by presenting the steps of our proposed estimation method, which contains cross-fitting (Chernozhukov et al.,, 2018) for separating the samples used for estimating the nuisance functions from that used for estimating the parameter of interest.
Estimating the nuisance functions:
We partition the samples into $L$ equal size folds $\{I_{1},\dots,I_{L}\}$. For $\ell\in\{1,\dots,L\}$, we estimate the nuisance functions $(\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})$ on data from all folds but $I_{\ell}$, which we denote as $I_{\ell}^{c}$. In order for estimating the nuisance functions, we proceed in the following two stages.
@@
      We choose $\hat{\pi}_{1,\ell}$ such that it satisfies
      $\mathbb{E}\left[\left\{(1-A)\hat{\pi}_{1,\ell}(X)-1\right\}h_{1}(X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{1}\in\mathcal{H}_{1},$ (1)
      where $\mathcal{H}_{1}\subset L^{2}(P_{0,X})$ is a function space chosen by the researcher.
    • We estimate $\mu_{1}$ by regressing $Y$ on $\{M,X,A=1\}$ to obtain $\hat{\mu}_{1,\ell}$.
@@
      Given $\hat{\pi}_{1,\ell}$ from Stage 1, we choose $\hat{\pi}_{2,\ell}$ such that it satisfies
      $\mathbb{E}\left[\left\{A\hat{\pi}_{2,\ell}(M,X)-(1-A)\hat{\pi}_{1,\ell}(X)\right\}h_{2}(M,X)\right]\approx 0,~{}~{}~{}~{}~{}~{}\forall h_{2}\in\mathcal{H}_{2},$ (2)
      where $\mathcal{H}_{2}\subset L^{2}(P_{0,(M,X)})$ is a function space chosen by the researcher.
    • Given $\hat{\mu}_{1,\ell}$ from Stage 1, we estimate $\mu_{2}$ by regressing $\hat{\mu}_{1,\ell}$ on $\{X,A=0\}$ to obtain $\hat{\mu}_{2,\ell}$.
Estimating the parameter of interest:
For any estimates of nuisance functions $\{\tilde{\pi}_{1},\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}\}$, let
$\phi(O;\tilde{\pi}_{1}$ $,\tilde{\pi}_{2},\tilde{\mu}_{1},\tilde{\mu}_{2}):=\tilde{\mu}_{2}(X)+(1-A)\tilde{\pi}_{1}(X)\{\tilde{\mu}_{1}(M,X)-\tilde{\mu}_{2}(X)\}+A\tilde{\pi}_{2}(M,X)\{Y-\tilde{\mu}_{1}(M,X)\}.$
For all $\ell$, let $\hat{\psi}^{mr}_{\ell}$ be the estimator of $\psi_{0}$ defined as,
$\hat{\psi}^{mr}_{\ell}=\mathbb{E}_{m}\left[\phi(O;\hat{\pi}_{1,\ell},\hat{\pi}_{2,\ell},\hat{\mu}_{1,\ell},\hat{\mu}_{2,\ell})\right],$
where the empirical expectation is evaluated on $I_{\ell}$. Our final estimator of the parameter of interest will be the following.
$\hat{\psi}^{mr}=\frac{1}{L}\sum_{\ell=1}^{L}\hat{\psi}^{mr}_{\ell}.$
The use of the cross-fitting approach plays a key role in establishing consistency and asymptotic normality of our proposed estimator, shown in Theorem 2.
We propose a non-parametric estimator for Stage 1 based on minimax optimization. A naive approach to implement this would be by solving the following optimization problem.
$\tilde{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)\right],$
where $\Pi_{1}\subset L^{2}(P_{0,X})$ is a function space chosen by the researcher. Unfortunately, this estimator will be unstable and highly sensitive to noise. The following result suggests an alternative.
@@
$\breve{\pi}_{1}$ obtained from the following minimax optimization satisfies Equation (1) with equality.
$\breve{\pi}_{1}=\arg\min_{\pi_{1}\in L^{2}(P_{0,X})}\max_{h_{1}\in L^{2}(P_{0,X})}\mathbb{E}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right].$
@@
$\hat{\pi}_{1}=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]+R_{n}^{1}(\pi_{1},h_{1}),$ (3)
where $R_{n}^{1}(\cdot,\cdot)$ is a regularizer. We will give a specific choice for $R_{n}^{1}(\cdot,\cdot)$ in Section 4.1.
Remark 1.
@@
Our proposed nonparametric estimator $\hat{\pi}_{1}$ in Equation (3) can also be used in the setting of Imai and Ratkovic, (2014) for estimating average treatment effect. Therefore, this provides a nonparametric version of the their covariate balancing propensity score (CBPS) method.
We can have similar arguments regarding $\pi_{2}$ in Stage 2.
Proposition 3.
For any give $\hat{\pi}_{1}$, $\breve{\pi}_{2}$ obtained from the following minimax optimization satisfies Equation (2) with equality.
$\breve{\pi}_{2}=\arg\min_{\pi_{2}\in L^{2}(P_{0,(M,X)})}\max_{h_{2}\in L^{2}(P_{0,(M,X)})}\mathbb{E}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right].$
Therefore, in Stage 2, to estimate $\pi_{2}$, we can use the following estimator.
$\hat{\pi}_{2}=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{A\pi_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]+R_{n}^{2}(\pi_{2},h_{2}),$ (4)
where $\Pi_{2}\subset L^{2}(P_{0,(M,X)})$ is a user-specified function class, and $R_{n}^{2}(\cdot,\cdot)$ is a regularizer.
@@
In order to estimate $\mu_{1}$ in Stage 1 and $\mu_{2}$ in Stage 2 of the approach, any standard parametric or non-parametric regression method can be used.
4.1 Doubly Robust Kernel Estimator
In this section, we describe solving the minimax problems (3) and (4) when RKHSes are used as the hypothesis classes. We use the method adopted by Ghassami et al., (2022), which builds upon the non-parametric estimation approach proposed by Dikkala et al., (2020) based on adversarial learning for solving integral equations within the semi-parametric proximal causal inference framework.
Let $\Pi_{1}$, $\mathcal{H}_{1}$, $\Pi_{2}$, and $\mathcal{H}_{2}$ be RKHSes with kernels $K_{\Pi_{1}}$, $K_{\mathcal{H}_{1}}$, $K_{\Pi_{2}}$ and $K_{\mathcal{H}_{2}}$, respectively, equipped with the RKHS norms $\|\cdot\|_{\Pi_{1}}$, $\|\cdot\|_{\mathcal{H}_{1}}$, $\|\cdot\|_{\Pi_{2}}$ and $\|\cdot\|_{\mathcal{H}_{2}}$. We propose the following Tikhonov regularization-based optimizations:
$\hat{\pi}_{1}$ $=\arg\min_{\pi_{1}\in\Pi_{1}}\max_{h_{1}\in\mathcal{H}_{1}}\mathbb{E}_{n}\left[\left\{(1-A)\pi_{1}(X)-1\right\}h_{1}(X)-\frac{1}{4}h^{2}_{1}(X)\right]-\lambda_{\mathcal{H}_{1}}\|h_{1}\|_{\mathcal{H}_{1}}^{2}+\lambda_{\Pi_{1}}\|\pi_{1}\|_{\Pi_{1}}^{2},$ (5)
$\hat{\pi}_{2}$ $=\arg\min_{\pi_{2}\in\Pi_{2}}\max_{h_{2}\in\mathcal{H}_{2}}\mathbb{E}_{n}\left[\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}h_{2}(M,X)-\frac{1}{4}h^{2}_{2}(M,X)\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\lambda_{\mathcal{H}_{2}}\|h_{2}\|_{\mathcal{H}_{2}}^{2}+\lambda_{\Pi_{2}}\|\pi_{2}\|_{\Pi_{2}}^{2},$ (6)
where $\lambda_{\mathcal{H}_{1}}$, $\lambda_{\Pi_{1}}$, $\lambda_{\mathcal{H}_{2}}$, and $\lambda_{\Pi_{2}}$ are regularization hyperparameters that are decreasing in $n$.
For the convergence analysis of the proposed minimax estimators, we direct the reader to (Dikkala et al.,, 2020; Ghassami et al.,, 2022).
Define $K_{\Pi_{1},n}=(K_{\Pi_{1}}(X_{i},X_{j}))_{i,j=1}^{n}$, $K_{\Pi_{2},n}=(K_{\Pi_{2}}((M,X)_{i},(M,X)_{j}))_{i,j=1}^{n}$, $K_{\mathcal{H}_{1},n}=(K_{\mathcal{H}_{1}}(X_{i},X_{j}))_{i,j=1}^{n}$, and $K_{\mathcal{H}_{2},n}=(K_{\mathcal{H}_{2}}((M,X)_{i},(M,X)_{j}))_{i,j=1}^{n}$ as the empirical kernel matrices corresponding to spaces $\Pi_{1}$, $\Pi_{2}$, $\mathcal{H}_{1}$, and $\mathcal{H}_{2}$, respectively. The following result provides closed-form solutions for $\hat{\pi}_{1}$ and $\hat{\pi}_{2}$.
Proposition 4.
Equation (5) achieves its optimum at $\hat{\pi}_{1}=\sum_{i=1}^{n}\alpha_{i}K_{\Pi_{1}}(X_{i},\cdot)$, with $\alpha$ defined as,
$\alpha=-\Big{(}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}diag(1-A)K_{\Pi_{1},n}+n^{2}\lambda_{\Pi_{1}}K_{\Pi_{1},n}\Big{)}^{\dagger}K_{\Pi_{1},n}diag(1-A)\Gamma_{1}e_{n},$
where $\Gamma_{1}=K_{\mathcal{H}_{1},n}(\frac{1}{n}K_{\mathcal{H}_{1},n}+\lambda_{\mathcal{H}_{1}}I_{n})^{-1}$, $diag(1-A)$ is a diagonal matrix with $1-A_{i}$ as the $i$-th diagonal entry, and $e_{n}:=\begin{pmatrix}1&1&\cdots&1\end{pmatrix}^{\top}$, and $\dagger$ denotes Moore-Penrose pseudo-inverse.
Equation (6) achieves its optimum at $\hat{\pi}_{2}=\sum_{i=1}^{n}\beta_{i}K_{\Pi_{2}}((M,X)_{i},\cdot)$, with $\beta$ defined as,
$\beta=-\Big{(}K_{\Pi_{2},n}diag(A)\Gamma_{2}diag(A)K_{\Pi_{2},n}+n^{2}\lambda_{\Pi_{2}}K_{\Pi_{2},n}\Big{)}^{\dagger}K_{\Pi_{2},n}diag(A)\Gamma_{2}\big{(}(1-A)\hat{\pi}_{1}(X)\big{)}_{n},$
where $\Gamma_{2}=K_{\mathcal{H}_{2},n}(\frac{1}{n}K_{\mathcal{H}_{2},n}+\lambda_{\mathcal{H}_{2}}I_{n})^{-1}$, $diag(A)$ is a diagonal matrix with $A_{i}$ as the $i$-th diagonal entry, and $\big{(}(1-A)\hat{\pi}_{1}(X)\big{)}_{n}:=\begin{pmatrix}(1-A_{1})\hat{\pi}_{1}(X_{1})&(1-A_{2})\hat{\pi}_{1}(X_{2})&\cdots&(1-A_{n})\hat{\pi}_{1}(X_{n})\end{pmatrix}^{\top}$.
5 Theoretical Analysis